1. What mass of glucose should you use to prepare 250.0 mL of a solution 0.112 M C6H12O6?
Answer #1

molarity of C6H12O6    = 0.112M

volume of solution        = 250ml   = 0.25L

no of moles of C6H12O6   = molarity * volume in L

                                           = 0.112*0.25   = 0.028 moles

mass of C6H12O6      = no of moles * gram molar mass

                                    = 0.028*180   = 5.04g >>>>answer